### v1

# Embedding-based statistical inference on generative models

Abstract
The recent cohort of publicly available generative models can produce human expert level content across a variety of topics and domains. Given a model in this cohort as a base model, methods such as parameter efficient fine-tuning, in-context learning, and constrained decoding have further increased generative capabilities and improved both computational and data efficiency. Entire collections of derivative models have emerged as a byproduct of these methods and each of these models has a set of associated covariates such as a score on a benchmark, an indicator for if the model has (or had) access to sensitive information, etc. that may or may not be available to the user. For some model-level covariates, it is possible to use “similar” models to predict an unknown covariate. In this paper we extend recent results related to embedding-based representations of generative models – the data kernel perspective space – to classical statistical inference settings. We demonstrate that using the perspective space as the basis of a notion of “similar” is effective for multiple model-level inference tasks.
corresponding author
1 Introduction
Generative models have recently met or surpassed human-level standards on benchmarks across a range of tasks (Nori et al., 2023; Katz et al., 2024; Dubey et al., 2024). While these claims warrant skepticism and further robustness evaluation (Ness et al., 2024), the impressive capabilities have created a competitive environment for training state-of-the-art models and have inspired the development of complementary methods to adapt models to particular use cases. For example, quantizing the weights of a neural model enables a trade-off of model precision with vRAM and disk space (Gholami et al., 2022) while methods such as Low Rank Adaptation (LoRA) (Hu et al., 2021) and prompt-tuning (Lester et al., 2021) enable compute- and data-efficient model adaptation. Other methods such as retrieval-augmented generation (Lewis et al., 2020), model merging (Matena & Raffel, 2022), constrained decoding (Hokamp & Liu, 2017), etc. (Dettmers et al., 2024; Edge et al., 2024), have similarly contributed to the rapid development of a large population of diverse and accessible models.
Each model in the population has an accompanying set of covariates – scores on benchmarks, training mixture proportions, model safety scores, probability of hallucination, etc. – that are a function of the model, the training set, the architecture, the retrieval database, or derivatives thereof. For a given model, the covariate of interest or the function to calculate it may not be available to the user. For example, a proprietary model may have been trained on copyrighted data and the indicator for whether or not the model has had access to the data is unknown. Methods for predicting model-level covariates are necessary in these settings, and others, to fully understand the behavior and properties of a model.
In this paper we extend recent theoretical results for embedding-based representations of generative models (Acharyya et al., 2024) to statistical inference settings. In particular, our results show that the embedding-based representations of a collection of models can be used for consistent inference for a wide class of inference problems. We demonstrate the effectiveness of the representations for three downstream inference tasks, including predicting the presence of sensitive information in a model’s training mixture and predicting a model’s safety. We include empirical investigations of performance sensitivity to hyperparameters required to induce the model representations.
1.1 Background & related work
Our work is an extension of recent theoretical and empirical results on embedding-based representations of generative models (Acharyya et al., 2024), which itself is a continuation of a long-line of embedding-based investigations of the inputs and outputs of generative models (Mikolov, 2013; Reimers, 2019; Neelakantan et al., 2022; Patil et al., 2023). Of particular relevance is recent empirical work defining the data kernel (Duderstadt et al., 2023) and investigations into its ability to track the dynamics of interacting models (Helm et al., 2024), in which the experiments demonstrate the ability to parlay the embeddings of a collection of inputs or outputs into useful vector representations of the generative models themselves in both white-box and black-box settings. The results herein further theoretically and empirically validate these findings in the context of statistical inference.
Our work is also related to using embedding-based techniques for inference on complicated objects such as entire mouse connectomes (Wang et al., 2020), physiological data (Chen et al., 2022), and classification distributions (Helm et al., 2021). In each of these settings, the authors define a pairwise distance matrix on the objects and apply multi-dimensionsal scaling to obtain vector representations of each of the objects. Once there is one vector representation per object, standard inference methods for the specific task can be used. The method we study herein follows this general formula, with the additional complication that the objects are random mappings.
Lastly, our work is a part of the relatively new literature on inference on generative models. For example, FlashHELM (Perlitz et al., 2023) uses the score on a subset of a benchmark such as HELM (Liang et al., 2022) to quickly identify where a model fits on a leaderboard. More recent work proposes scaling laws to predict the performance of base models on a suite of benchmarks as a function of training FLOPs (Ruan et al., 2024). The method proposed herein does not assume access to a scoring function nor does it assume access to a function of the model’s weights at inference time. Perhaps most importantly, our setting and method can be applied to general model-level prediction scenarios.
2 The Data Kernel Perspective Space
Following Acharyya et al. (2024), we consider a generative model $f$ with weights $W$ and decoding function (temperature, maximum response length, etc.) $\delta$ to be a random map from a query space $\mathscr{Q}$ to a response space $\mathcal{X}$. We let $f_{1},\ldots,f_{n}$ be $n$ such random maps for a shared $\mathscr{Q}$ and $\mathcal{X}$. For example, $f_{i}$ may be the neural model parameterized by the fixed context augmentation $a_{i}$, the retrieval vector database $V_{i}$, the fine-tuned model with weights $W+\Delta W_{i}$ or the model with decoding function $\delta_{i}$. Given a collection of queries $Q=\{q_{1},\ldots,q_{m}\}$ we let $F_{ij}$ denote the distribution on the response space for fixed $f_{i}$ and $q_{j}$ and assume model responses $f_{i}(q_{j})_{1},\ldots,f_{i}(q_{j})_{r}=x_{ij1},\ldots,x_{ijr}$ are $i.i.d.$ realizations from $F_{ij}$.
Our goal is to induce a representation of the neural models that captures information relevant to model-level statistical inference. We abuse notation by letting $x$ represent a realization of a random variable and the random variable itself and include necessary context when there is potential for confusion.
We let $g:\mathcal{X}\to\mathbb{R}^{p}$ be an embedding function that maps a response to a $p$-dimensional real-valued vector and, for notational convenience, let $\bar{x}_{ij}=\frac{1}{r}\sum_{k=1}^{r}g(x_{ijk})$ be the average of the embedded responses. We use $\bar{X}_{i}\in\mathbb{R}^{m\times p}$ to denote the matrix where the $j$th row is $\bar{x}_{ij}$ and view this matrix as a representation of model $f_{i}$ with respect to $Q$ and $g$.
We define $D$ as the $n\times n$ pairwise distance matrix with entries
$D_{ii^{\prime}}=\frac{1}{m}\big{|}\big{|}\bar{X}_{i}-\bar{X}_{i^{\prime}}\big{|}\big{|}_{F}.$ (1)
Each entry $D_{ii^{\prime}}$ is the average difference between the average embedded response between model $f_{i}$ and model $f_{i^{\prime}}$ and captures the difference in average model behavior with respect to $Q$ and $g$. With the form described by Eq. (1), $D$ is a Euclidean distance matrix and the multidimensional scaling (MDS) of $D$ yields $d$-dimensional Euclidean representations of the matrices $\bar{X}_{i}$ (Torgerson, 1952) and, thus, Euclidean representations of the models $f_{1},\ldots,f_{n}$ with respect to $Q$ and $g$. Letting $\widehat{\psi}:=\text{MDS}(D)$, we refer to the $i$th row of $\widehat{\psi}_{i}\in\mathbb{R}^{d}$ as the perspective of model $f_{i}$ and the entire configuration $\widehat{\psi}$ as the data kernel perspective space (DKPS).
2.1 Analytical properties of the data kernel perspective space
While the perspectives are Euclidean objects, it is not possible to comment on their properties without imposing constraints on the $F_{ij}$. To facilitate analysis we assume each query is from the query distribution $G$ with support $\mathscr{Q}$ and let $\mu_{i}\in\mathbb{R}^{m\times p}$ denote the matrix whose $i^{th}$ row is $\mathbb{E}_{x\sim F_{ij}}\left[g(x)\right]$. We let $\Delta$ be such that $\Delta_{ii^{\prime}}=\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$. For settings where $m$ grows and $q_{j}\overset{iid}{\sim}G$, we refer to $\mu_{i}$ as $\mu_{i}(G)$ and let $\Delta^{*}(G)$ be such that $\Delta^{*}_{ii^{\prime}}(G)=\lim_{m\to\infty}\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$. Finally, we assume $g$ is bounded. [^1] Boundedness is typically satisfied in practice for well-defined $g$: for language models, an element of $\mathcal{X}$ is a finite sequence from a finite vocabulary; for text-to-image models $\mathcal{X}=\{0,\ldots,255\}^{3}$ is finite.
With $q_{1},\ldots,q_{m}\overset{iid}{\sim}G$, and $x_{ij1},\ldots,x_{ijr}\overset{iid}{\sim}F_{ij}$, we have
$D_{ii^{\prime}}=\frac{1}{m}\left\lVert\bar{X}_{i}-\bar{X}_{i^{\prime}}\right\rVert\to\Delta^{*}_{ii^{\prime}}(G)$ (2)
with high probability as $m,r\to\infty$ for all $(i,i^{\prime})\in\{1,\ldots,n\}\times\{1,\ldots,n\}$. Under technical assumptions described in Acharyya et al. (2024), there exists $\psi(G):=\text{MDS}(\Delta^{*}(G))$ such that $\widehat{\psi}\to\psi(G)$. The configuration $\psi(G)$ captures the true geometry of the mean descrepancies of the model responses with respect to queries from $G$. We drop the argument of $\psi$ when the distribution of the queries is clear. In a more general setting where the models are assumed to be $i.i.d.$ realizations from a model distribution $F_{model}$, the distance matrix $D$ converges to $\Delta^{*}$ and, under technical assumptions, there exists $\psi$ such that $\widehat{\psi}\to\psi$ as $m,r\to\infty$, for all $n$ (Acharyya et al., 2024).
In settings where the query space contains only a single element or where each $F_{ij}$ is a point mass on $x_{ij}$, $\Delta_{ii^{\prime}}$ is exactly the maximum mean discrepancy (Gretton et al., 2012) of the distributions $F_{ij}$ and $F_{i^{\prime}j}$ under a linear map applied to $g(x)$. For settings where $|\mathscr{Q}|>1$ and the $F_{ij}$ are not all point masses, the analysis is more complicated since both $F_{ij}$ and $\bar{x}_{ij}$ are random variables.
2.2 Statistical inference in the data kernel perspective space
Consider the classical statistical learning problem (Hastie et al., 2009, Chapter 2): Given training data $\mathcal{T}_{n}=\{(X_{1},y_{1}),\ldots,(X_{n},y_{n})\}$ assumed to be $i.i.d.$ realizations from the joint distribution $P_{XY}$, select the decision function $h:\mathcal{X}\to\mathcal{Y}$ that minimizes the expected value of a loss function $\ell$ with respect to $P_{XY}$ within a class of decision functions $\mathcal{H}$ for a test observation $X$ assumed to be an independent realization from the marginal distribution $P_{X}$. Or, with $\mathcal{R}_{\ell}(P_{XY},h):=\mathbb{E}_{P_{XY}}\left[\ell(h(X),y)\right]$, select $h$ such that
$h\in\operatornamewithlimits{argmin}_{h\in\mathcal{H}}\mathcal{R}_{\ell}(P_{XY},h).$
We let $\mathcal{R}_{\ell}^{*}(P_{XY},\mathcal{H})$ denote the expected loss of $h$ in the argmin and let $\mathcal{R}^{*}_{\ell}(P_{XY})$ denote the minimum $\mathcal{R}_{\ell}^{*}(P_{XY},\mathcal{H})$ over all $\mathcal{H}$.
The joint distribution $P_{XY}$ is often unavailable and the decision function is selected based on $\mathcal{T}_{n}$. We let $h(\;\cdot\;;\mathcal{T}_{n})$ denote such a decision function and say the sequence of decision functions $\left(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n})\right)$ is consistent for $P_{XY}$ with respect to $\mathcal{H}$ if
$Pr\big{(}\;\big{|}\mathcal{R}_{\ell}(P_{XY},h(\;\cdot\;;\mathcal{T}_{n}))-\mathcal{R}^{*}_{\ell}(P_{XY},\mathcal{H})\big{|}>\epsilon\big{)}\to 0$
as $n\to\infty$.
In our setting we observe $\widehat{\mathcal{T}}_{n}=\{(\widehat{\psi}_{1},y_{1}),\ldots,(\widehat{\psi}_{n},y_{n})\}$ and the test observation $\widehat{\psi}_{n+1}$ where $\widehat{\psi}_{i}$ is an errorful observation of the true-but-unknown $\psi_{i}$ and $y_{i}$ is the model-level covariate corresponding to $f_{i}$. Two important extensions of the convergence results described in (Acharyya et al., 2024) are the following theorems:
Theorem 1.
Under technical assumptions described in Appendix A,
$\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))\to\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\mathcal{T}_{n}))$
as $m,r\to\infty$;
and
Theorem 2.
Under technical assumptions described in Appendix A, if $(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$ then $(h(\;\cdot\;;\widehat{\mathcal{T}}_{1}),\ldots,h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$ as $n,m,r\to\infty$.
The proofs of Theorems 1 and 2 are provided in Appendix A.
These results do not say that with enough models, queries, and responses that inference using the $\widehat{\psi}$ is “optimal”. Instead, Theorem 1 says that with enough queries and responses for each query that the performance of a decision function that uses $\widehat{\mathcal{T}}_{n}$ will be “close” to the performance of the decision function that uses $\mathcal{T}_{n}$ – regardless of how far away $\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\mathcal{T}_{n}))$ is from $\mathcal{R}_{\ell}^{*}(P_{\psi Y},\mathcal{H})$.
Similarly, Theorem 2 says that if there is a sequence of decision functions known to be consistent with respect to $\mathcal{H}$ when using $\mathcal{T}_{n}$, then the same sequence of decision functions is consistent with respect to $\mathcal{H}$ when using $\widehat{\mathcal{T}}_{n}$ where $m$ and $r$ are allowed to grow. The result does not provide instructions on how to construct such a sequence. Nor does it provide insight as to how close $\mathcal{R}_{\ell}^{*}(P_{\psi Y},\mathcal{H})$ is to $\mathcal{R}_{\ell}^{*}(P_{\psi Y})$. Most importantly – and most unique to our setting – Theorem 2 does not provide guidance for how to select a query distribution $G$ such that $|\mathcal{R}_{\ell}^{*}(P_{fY})-\mathcal{R}_{\ell}^{*}(P_{\psi Y})|$ is minimal.
While we do not provide theoretical details as to which situations optimal model performance is possible, if $y_{i}$ can be described as a function of the average of model responses for a set of queries from an optimal $G^{*}$ then $\psi(G^{*})$ will be such that $\mathcal{R}_{\ell}^{*}(P_{fY})=\mathcal{R}_{\ell}^{*}(P_{\psi Y})$ (Devroye et al., 2013, Chapter 32). Conversely, it is unclear how effective decision functions constructed with $\psi(G^{*})$ for model-level covariates that cannot be described as a function of the average response will behave. Further, it is unclear how “close” a query distribution $G$ needs to be to $G^{*}$ to induce an informative representation of the models with respect to $y_{i}$. We demonstrate a potential trade-off between the proximity of $G$ to $G^{*}$ and the number of queries necessary to tease out model differences in one of the experimental settings below.
Lastly, we note that our setting, methodology, and theorems are general to classes of generative models in which there exists a well-defined embedding function from the output space to $\mathbb{R}^{p}$, though our experimental settings focus entirely on collections of language models.
2.3 An illustrative example – “Was RA Fisher great?”
Figure 1: Left. The 2-d Data Kernel Perspective Space (DKPS) and covariate surface for a collection of 550 models parameterized by fixed augmentations. Right. The performance of the 1-nearest neighbor regressor in DKPS for predicting the probability that an unlabeled model responds “yes” to “Was RA Fisher great?”.
The first model-level inference task we study is a toy example where the task is to predict the probability that a model will respond “yes” to the question “Was RA Fisher great?”. The question is chosen due to its subjectiveness – there is no correct answer to the question of a person’s greatness – and its duality – Ronald A. Fisher is considered one of the most influential statisticians in history and is considered an advocate of the 20th century Eugenics movement (Bodmer et al., 2021).
We use a 4-bit version of Meta’s LLaMA-2-7B-Chat (Touvron et al., 2023) as a base model and consider a collection of models parameterized by fixed context augmentations. Each augmentation $a_{i}$ contains information related to RA Fisher’s statistical achievements (i.e., $a_{i}$ = “RA Fisher pioneered the principles of the design of experiments”) or to his involvement in the eugenics movement (i.e., $a_{i}$ = “RA Fisher’s view on eugenics were primarily based on anecdotes and prejudice.”) and is prepended to every query. The covariate corresponding to a given model is calculated by prompting the base model with the appropriately formatted prompt “Give a precise answer to the question based on the context. Don’t be verbose. The answer should be either a yes or a no. $a_{i}$. Was RA Fisher great?” until there are 100 valid responses. We let $y_{i}$ be the average number of “yes”es.
To induce a DKPS for this task, we consider queries sampled from OpenAI’s ChatGPT with the prompt “Provide 100 questions related to RA Fisher.”. For a given query $q_{j}$ we prompt the base model with the appropriately formatted prompt “$a_{i}\;q_{j}$” and fix $r=1$. The left figure of Figure 1 is a 3-d figure where the first two dimensions are the DKPS of the $n=550$ ($275$ statistics augmentations, $275$ eugenics augmentations) models induced with $m=100$ queries. Model responses are embedded by averaging the per-token last layer activation of the base model. The third dimension is an interpolated $y_{i}$ surface with a linear kernel (Du Toit, 2008). The first dimension of the DKPS is clearly capable of distinguishing between models adorned with “statistics” augmentations and models adorned with “eugenics” augmentations. The shape of the interpolated covariate surface is highly correlated with this feature. A description of the augmentations that parameterize the models and the queries used to induce the DKPS is provided in Appendix B.1.
The right figure of Figure 1 shows the performance of a 1-nearest neighbor (1-NN) regressor in DKPS for a varying number of labeled models and a varying number of queries. The DKPS is induced with $n$ models and the regressor is trained with $n-1$ of the model-level covariates. The mean squared error reported is the error of the 1-NN regressor for predicting the “left out” model’s covariate ($\pm$ 1 S.E.). As Theorems 1 and 2 suggest, the performance of the regressor is dependent on both the number of models and the number of queries: the more models and the more queries the better. The scale of the impact of more models and more queries, however, depends on its counterpart. The amount of models does not have a large impact on predictive performance if the number of queries is small. The amount of queries has a large impact on performance regardless of the number of models.
3 Experiments
We next consider two experiments with more realistic model-level covariates: predicting whether or not a model has had access to sensitive information and predicting model safety. For all experiments we fix $r=1$ as suggested by the empirical rates of convergence of the perspectives described in Acharyya et al. (2024). We use the MDS implementation from Graspologic (Chung et al., 2019) throughout. We use the profile likelihood of the singular values of $D$ to determine the dimensionality of the DKPS (Zhu & Ghodsi, 2006) and note that this may be larger than two. We show only the first two dimensions for visualization purposes.
3.1 Has a model seen sensitive information?
Figure 2: Left. The 2-d data kernel perspective space (DKPS) of 50 fine-tuned models – 25 with “sensitive” data in the fine-tuning data mixture (red), 25 with none (black) – induced by an evaluation set containing 10 prompts relevant to the sensitive data. For models trained on sensitive data, color intensity correlates with amount of sensitive data in the training mixture. Center. The 2-d DKPS of the models induced by a set of 10 prompts “orthogonal” to the difference between models with sensitive data in their fine-tuning data mixture and models with no sensitive data in their fine-tuning data mixture. Right. Classification performance as a function of number of labeled models and size of evaluation set for both sensitive and orthogonal evaluation sets.
Modern language models are trained with trillions of tokens of text (Touvron et al., 2023). For proprietary models such as OpenAI’s GPT series or Anthropic’s Claude series, the exact sources of the training mixtures are unknown and – given the models’ propensities to produce content that is strikingly similar to copyrighted content (Henderson et al., 2023) – its curation and use is ethically questionable (Lemley & Casey, 2020). Further, the training mixture of some models may include sensitive informative such as personal information, trade secrets, or government-classified information that should never be presented to the end user. Developing classifiers to identify models that are either trained on sensitive or copyrighted information or are likely to produce sensitive or copyrighted information is thus paramount to uphold the rights of the stakeholders of the original content.
To investigate the utility of DKPS for this purpose, we again use a 4-bit version of LLaMA-2-7B-Chat as a base model and train 50 different LoRA adapters with different subsets of the Yahoo! Answers (YA) dataset (Zhang et al., 2015). The YA dataset consists of data from 10 topics. We consider data from the topic “Politics & Government” to be “sensitive” information, data from the topics “Society & Culture”, “Science & Mathematics”, “Health”, “Education & Reference”, “Computers & Internet”, and “Sports” to be “not-sensitive”, and data from the remaining topics to be “orthogonal”. We trained each of the 50 adapters with 500 question-answer pairs for 3 epochs with a learning rate of $5\times 10^{-5}$ and a batch size of 8. Each adapter is rank 8, has a scaling factor of 32, targets all attention layers, does not have bias terms, and has a dropout probability of 0.05 when training. For 25 of the adapters, the adapter training mixture consisted wholly of randomly selected not-sensitive data. For the remaining 25 adapters, the adapter training mixture consisted of $p_{i}$ randomly selected sensitive data and $500-p_{i}$ randomly selected not-sensitive data. We let $y_{i}$ be the indicator of whether or not the adapter’s training mixture contained any sensitive data.
We study classification of the models in two different DKPS: one induced by a set of randomly selected sensitive queries, one induced by a set of randomly selected orthogonal queries. Both DKPS use the open source embedding model nomic-embed-v1.5 (Nussbaum et al., 2024). A 2-d DKPS induced by $m=10$ randomly selected sensitive queries is shown on the left of Figure 2. In this space the models are separated by their label and the models that have seen more sensitive information are generally farther from the class-boundary. A 2-d DKPS induced by $m=10$ orthogonal queries is shown in the center of Figure 2. Here, by contrast, the models are not easily separable by their label.
The right figure of Figure 2 shows the classification performance of Fisher’s Linear Discriminant trained on varying amounts of DKPS representations of models for the two query distributions. For a given $n$ and $m$, we induce the DKPS with all
Figure 3: Top. The 1-d FLD projection of the models from a DKPS induced by queries from the sensitive topic versus the amount of sensitive data the adapter had access to during training. Bottom. The same but for a DKPS induced by queries from orthogonal topics.
models and a randomly selected query set. We report the expected risk of the classifier trained on a random subset of the models for the remaining models. We observe similar phenomena to the “Was RA Fisher great?” experiment in that both the amount of models and the amount of queries impact performance. For a fixed $m$, the expected risk curves also highlight the observed difference in separability of the models in the two DKPS, with the expected risk with sensitive queries being significantly lower than the expected risk with orthogonal queries. Indeed, the expected risk with $m=10$ sensitive queries is similar to the expected risk with $m=50$ orthogonal queries.
Lastly, we highlight the 1-d projection learned by Fisher’s linear discriminant for $m=10$ for both DKPS in Figure 3. As can be seen in the top figure of Figure 3, the projection of the models is correlated with the proportion of sensitive data that the model has had access to when the DKPS is induced with sensitive queries. While the linear goodness-of-fit is not large ($R^{2}=0.37$), the correlation is statistically significant per the hypothesis test using Kendall’s rank correlation coefficient ($\tau=0.42,p<0.01$). The projection of the models when the DKPS is induced with orthogonal queries has a line of best fit with a negligible slope and a much smaller Kendall’s rank correlation coefficient ($\tau=0.08$). The second of which results in a $p$ value of $0.57$ – meaning we fail to reject the null that the amount of sensitive information the model has had access to is correlated with the learned 1-d projection.
Throughout this experiment we use the term “orthogonal” for queries from topics that are not used when fine-tuning the models under study. The term is chosen because, naïvely, queries from these topics should not elicit different responses from models trained on sensitive data and models trained on not-sensitive data. In reality, the sensitive data and the not-sensitive data have different underlying token distributions and this difference will cause systematic differences in model responses after fine-tuning even for topics that are irrelevant a priori. Further, it is likely that the documents in the “orthogonal” topics share some content commonalities with documents in the sensitive and not-sensitive topics. We see this phenomenon in the classification results in the right figure of Figure 2 where the linear classifier is able to perform better than chance with enough “orthogonal” queries.
3.2 How safe is a model?
Figure 4: Left. A graph where each node is a model and an edge between two models exists if model $i$ is fine-tuned from model $i^{\prime}$ or if model $i^{\prime}$s weights were used in a model-merge that resulted in model $i$, etc. Right, top. The two-dimensional data kernel perspective spaces (DKPS) corresponding to the toxicity and bias prediction tasks. Dot size is proportional to model toxicity or bias. Right, bottom. Relative performance of three regression techniques. Local predictions in DKPS are more effective than both global predictions and local predictions in model relationship space.
Model safety is one of the biggest concerns when deploying a language model in production. An unsafe model is prone to propagating harmful stereotypes (Ferrara, 2024), using toxic language (Wen et al., 2023), and misunderstanding the user’s intent (Ji et al., 2023) – all of which can adversely affect the user and their experience. Hence, developing techniques to understand how unsafe a model is an important aspect of the model-production pipeline. As with predicting if a model has had access to sensitive information above, we investigate using DKPS to predict model safety through the lens of model toxicity and model bias.
We consider a collection of 58 models. Each model in the collection is a base model, a fine-tuned version of a base model, or the result of weight-merging various other models in the collection. We view this collection of models as a graph where each model is a node and an undirected edge exists between nodes if one of the models is a fine-tuned version of the other or if one of the models is the result of a model merge including the other. The graph representing the collection of models is shown on the left of Figure 4. The list of models under study is provided in Appendix B.2.
For each model in the collection we consider two covariates: model toxicity and model bias. To determine a model’s toxicity, we prompt each model with a collection of queries from the Real Toxicity Prompts (RTP)
Figure 5: The relative time improvement (larger is better) when using local predictions in DKPS instead of calculating the ground-truth model-level covariate using HuggingFace’s API.
(Gehman et al., 2020) dataset and subsequently evaluate the toxicity of each response with the neural model roberta-hate-speech-dynabench-r4 (Vidgen et al., 2021). The model-level toxicity is simply the average response toxicity. An analagous process is used to define a model’s bias with the dataset Bias in Open-ended Language Generation Dataset (BOLD) (Dhamala et al., 2021) and the regard model (Sheng et al., 2019).
We induce the perspective spaces for the toxicity and bias tasks with randomly sampled prompts from RTP and BOLD, respectively, and the embedding model nomic-embed-v1.5. The 2-d DKPS – induced with $m=2000$ queries – for both tasks is shown in the top right of Figure 4. The size of the dot is correlated with the model’s covariate. We highlight an example unlabeled model (green) and its corresponding neighbor in DKPS (red) and neighbors in graph-space (blue) in Figure 4. Importantly, the relative position of a model in the respective DKPS is predictive of the model’s toxicity and the model’s bias.
We quantify this observation by evaluating regressors for predicting the model-level toxicity and bias of an unlabeled model. We consider three different regressors for these tasks. The first is a constant equal to the average covariate of the labeled models (i.e., the “global mean”). The second uses the average covariate of models who share an edge with the unlabaled model (i.e., “1-NN (graph)”). The third uses the covariate of the nearest neighbor in DKPS (i.e., “1-NN (DKPS)”). For 1-NN (DKPS) we consider varying amounts of randomly sampled queries from RTP and BOLD to induce the DKPS. We use the global mean as a standard and report the relative absolute error of the three methods in the bottom right of Figure 4. The reported performance of the 1-NN (DKPS) regressor is the average of the smaller of $200$ and $2000/m$ random samples of $m$ queries. Notably, given enough queries local predictions in DKPS outperform local predictions in the graph space and predictions using the global mean.
In addition to reporting the relative performance, we report the time it takes to use the DKPS machinery to predict Mistralv1.0’s (Jiang et al., 2023) toxicity and bias relative to using the evaluation model through HuggingFace’s API (Wolf et al., 2019). The time we report for the HuggingFace API is the time it takes to calculate the “ground truth” used to calculate the performance of the regressors above and is the total time required for Mistralv1.0 to produce responses and for the evaluation model to produce scores for 2000 queries. The time we report for 1-NN (DKPS) includes the time it takes for Mistralv1.0 to produce $m$ responses, the time it takes nomic-embed-v1.5 to embed the responses, the time it takes to induce the DKPS, and the time it takes to train and use the nearest neighbor regressor. It does not include the time it takes to produce and evaluate the responses of the other models in the collection. The relative efficiency of DKPS, as seen in Figure 5, is approximately $1/m$. This relationship will hold for all model-level inference tasks where the covariate is proportional to the sum of a function of individual responses.
4 Discussion
We have demonstrated – both theoretically and empirically – that embedding-based representations of generative models can be used for various model-level inference tasks. While the results we presented show the potential of our approach, there are choices throughout the collection-of-models-to-covariate-prediction pipeline that can affect performance and implementation practicality.
As mentioned in Section 2.2, one major decision is the query distribution $G$ (or, more practically, the query set $Q$). In particular, the representations of the models that the query set induces may or may not be relevant to a particular inference task. We demonstrate this phenomenon in Section 3.1 where the queries from the “sensitive-only” query distribution can induce representations of similar discriminative ability as queries from the “orthogonal-only” query distribution with $1/5$ of the queries. We expect a similar but less dramatic effect within the distributions of “sensitive-only” queries and anticipate that curating an “optimal” set of queries for a given $g$ and for a fixed $|Q|$ will soon be a highly active research area.
Another decision is the distance function used to define $D$. The MDS of the distance matrix studied herein (Eq. (1)) produces representations of the models that are consistent for objects that capture the true mean discrepancy geometry of the model responses. For model-level covariates that cannot naturally be described as a function of mean discrepancy geometry, we do not expect that information-theoretic optimal performance when using $\widehat{\psi}(G)$ is possible for any $n,m,$ and $r$. Instead, for representations of the models that can be used generally without information loss, it is necessary to replace the Frobenius norm of the differences of the average embedded response with a more expressive distance such as an extension of a distance defined directly on the cumulative distributions of responses or with a task-specific distance (Helm et al., 2020). Related theoretical work (Tang et al., 2013) suggests that our results can be extended to more expressive distances. We do not expect the naïve replacement of the Frobenius norm with a more expressive distance function to be universally better for model-level inference tasks in practice as there are likely computational and query set quality trade-offs to consider. As with the active curation of an optimal query set, we expect these trade-offs to be important future research topics.
In the experiments above we have access to $n$ models and $n^{\prime}<n$ corresponding model-level covariates. We induce a DKPS with the entire collection of models and use the $n^{\prime}$ labeled models to predict a label for the remaining $n-n^{\prime}$ models. In practice an unlabeled model may not be available when inducing the DKPS or, if $n$ is large, it may be too expensive to induce a DKPS whenever a prediction for a new unlabeled model is required. Out-of-sample techniques (Bengio et al., 2003; Trosset & Priebe, 2008) can be used to meet these imposed constraints at the cost of a slight degradation in representation quality and, hence, inference performance.
Implementation of inference in DKPS in practice will require an upfront, one-time cost of generating responses from a subset of models in the collection and scoring their outputs. Since this is required to compare the models with respect to the covariates anyway we followed the timing paradigm presented in (Perlitz et al., 2023) and did not include this cost when comparing the methods in Figure 5. Once the models in the initial subset are scored we expect the relative time efficiency (and implied relative computational efficiency) of inference in DKPS to be worthwhile to practitioners.
We fixed $r=1$ and let $m$ grow throughout our experiments. For Theorems 1 and 2 to hold, both $m$ and $r$ must grow. In practice, the trade-off between getting responses for more queries or getting more responses for the same queries depends on the distributions $F_{ij}$. For example, if $F_{ij}$ is a point mass then $r>1$ for $q_{j}$ is unnecessary. Conversely, if $F_{ij}$ is a complicated distribution on $\mathbb{R}^{p}$ then more responses are necessary to properly estimate it and, hence, to properly capture the difference between average model responses.
Lastly, we note that the data kernel perspective space could be used to extend the statistical Turing test framework presented in Helm et al. (2023) by comparing the conditional distributions of populations of humans and populations of machines in low-dimensional Euclidean space.
Acknowledgements
We would like to thank Avanti Athreya, Tianyi Chen, Ben Johnson, Michael Trosset, Tim Wang, Zekun Wang, and Weiwei Yang for providing helpful comments throughout the development of this manuscript.
Appendix A Proofs of Theorems 1 & 2
We introduce some notation to make the proofs of Theorems 1 and 2 easier to read. Bold letters (such as $\mathbf{B}$ or $\boldsymbol{\mu}$) are used to represent vectors and matrices. Any vector by default is a column vector. For a matrix $\mathbf{B}$, the $j$-th row is denoted by $(\mathbf{B})_{j\cdot}$, and the $(i,i^{\prime})$-th entry is denoted by $\mathbf{B}_{ii^{\prime}}$. Moreover, $\left\lVert\mathbf{B}\right\rVert_{F}$ denotes the Frobenius norm of the matrix $\mathbf{B}$. For any two vectors $\mathbf{x}$ and $\mathbf{y}$, $\left\lVert\mathbf{x}-\mathbf{y}\right\rVert$ denotes the Euclidean distance between $\mathbf{x}$ and $\mathbf{y}$. The set $\{1,2,\dots n\}$ is denoted by $[n]$. The set of $d\times d$ orthogonal matrices is denoted by $\mathcal{O}(d)$. For a sequence of random variables $X_{1},\ldots,X_{n}$, we say $X_{n}$ converges in probability to $X$ if $\lim_{n\to\infty}Pr\left[\left\lVert X_{n}-X\right\rVert>\epsilon\right]=0$ for every $\epsilon>0$. We denote convergence in probability with $X_{n}\to^{P}X$.
Recall that our setting includes the observed training set $\widehat{\mathcal{T}}=\{(\widehat{\psi}_{1},y_{1}),\ldots,(\widehat{\psi}_{n},y_{n})\}$ with the true-but-not-observed $(\psi_{i},y_{i})\overset{iid}{\sim}P_{\psi Y}$ and realizations $\psi_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}^{d^{\prime}}$, a unlabeled test observation $\widehat{\psi}_{n+1}$, a class of decision functions $\mathcal{H}\subset\{h:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}\}$, and a loss function $\ell:\mathbb{R}^{d^{\prime}}\times\mathbb{R}^{d^{\prime}}\to\mathbb{R}_{\geq 0}$.
A.1 Proof of Theorem 1
Define
$\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))$ $=\mathbb{E}_{(\psi_{i},y_{i})_{i\in[n+1]}\overset{iid}{\sim}P_{\psi Y}}[l(h(\widehat{\psi}_{n+1};\{(\widehat{\psi}_{i},y_{i})\}_{i=1}^{n}),y_{n+1})]$
and, analogously,
$\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\mathcal{T}_{n}))$ $=\mathbb{E}_{(\psi_{i},y_{i})_{i\in[n+1]}\overset{iid}{\sim}P_{\psi Y}}[l(h(\psi_{n+1};\left\{(\psi_{i},y_{i})\right\}_{i=1}^{n}),y_{n+1})].$
Recall that $n$ remains fixed. Following Acharyya et al. (2024), we let $m$ grow with the number of replicates $r$. Thus, $\widehat{\psi}_{i}$ depends on $r$. We write $\widehat{\psi}_{i}^{(r)}$ to emphasize this dependence when necessary. Note that $\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))$ also depends on $r$ (and $m$, through $r$).
We make some assumptions about the decision function $h$ and the loss function $l$.
Assumption 1. The decision function $h$ is invariant to affine transformation. That is, for any $\mathbf{W}\in\mathcal{O}(d)$ and $\mathbf{a}\in\mathbb{R}^{d}$,
$h(\mathbf{W}\psi_{n+1}+\mathbf{a};\left\{(\mathbf{W}\psi_{i}+\mathbf{a},y_{i})\right\}_{i=1}^{n})=h(\psi_{n+1};\left\{(\psi_{i},y_{i})\right\}_{i=1}^{n}).$
Assumption 2. The decision function $h$ is continuous. That is, if
$\max_{i\in[n]}\left\lVert\widehat{\psi}_{i}^{(r)}-\psi_{i}\right\rVert\to 0\text{ as }r\to\infty,$
then
$\left\lVert h\left(\widehat{\psi}_{n+1}^{(r)};\{(\widehat{\psi}_{i}^{(r)},y_{i})\}_{i=1}^{n}\right)-h\left(\psi_{n+1};\left\{(\psi_{i},y_{i})\right\}_{i=1}^{n}\right)\right\rVert\to 0.$
Assumption 3. We assume that the $\mathcal{H}$ is such that for every $h\in\mathcal{H}$, the image set of the function $h$ is closed, bounded and complete.
Assumption 4. The loss function $l$ is continuous. That is, for every $y\in\mathbb{R}^{d^{\prime}}$, $\left\lVert l(h^{\prime},y)-l(h^{\prime\prime},y)\right\rVert\to 0$ if $\left\lVert h^{\prime}-h^{\prime\prime}\right\rVert\to 0$.
Thus, by Theorem 2 of Acharyya et al. (2024), we can say that there exist sequences $\{\mathbf{W}^{(u)}\}_{u=1}^{\infty}$ and $\{\mathbf{a}^{(u)}\}_{u=1}^{\infty}$, where $\mathbf{W}^{(u)}\in\mathcal{O}(d)$ and $\mathbf{a}^{(u)}\in\mathbb{R}^{d}$ for all $u\in\mathbb{N}$, such that
$\max_{i\in[n]}\left\lVert\widehat{\psi}_{i}^{(r_{u})}-(\mathbf{W}^{(u)}\psi_{i}+\mathbf{a}^{(u)})\right\rVert\to 0\text{ as $u\to\infty$}.$ (3)
Now,
$\mathcal{R}_{\ell}(P_{\psi Y},h(\;.\;;\hat{\mathcal{T}}_{n}))-\mathcal{R}_{\ell}(P_{\psi Y},h(\;.\;;\mathcal{T}_{n}))$
$=\mathbb{E}\left[l(h(\widehat{\psi}_{n+1}^{(r_{u})};\{(\widehat{\psi}_{i}^{(r_{u})},y_{i})\}_{i=1}^{n}),y_{n+1})-l(h(\psi_{n+1};\left\{(\psi_{i},y_{i})\right\}_{i=1}^{n}),y_{n+1})\right]$
$=\mathbb{E}\left[l(h(\widehat{\psi}_{n+1}^{(r_{u})};\{(\widehat{\psi}_{i}^{(r_{u})},y_{i})\}_{i=1}^{n}),y_{n+1})-l(h(\mathbf{W}^{(u)}\psi_{n+1}+\mathbf{a}^{(u)};\{(\mathbf{W}^{(u)}\psi_{i}+\mathbf{a}^{(u)},y_{i})\}_{i=1}^{n}),y_{n+1})\right]$
from Assumption 1.
Using Assumption 2 on Eq. 3, we have
$\left\lVert h(\widehat{\psi}_{n+1}^{(r_{u})};\{(\widehat{\psi}_{i}^{(r_{u})},y_{i})\}_{i=1}^{n})-h\left(\psi_{n+1};\left\{(\psi_{i},y_{i})\right\}_{i=1}^{n}\right)\right\rVert\to^{P}0\text{ as $u\to\infty$}.$
Further, using Assumption 4, we get
$\left|l(h(\widehat{\psi}_{n+1}^{(r_{u})};\{(\widehat{\psi}_{i}^{(r_{u})},y_{i})\}_{i=1}^{n}),y_{n+1})-l(h(\psi_{n+1};\{(\psi_{i},y_{i})\}_{i=1}^{n}),y_{n+1})\right|\to^{P}0\text{ as $u\to\infty$,}$
which leads us to
$\left|\mathcal{R}_{\ell}(P_{\psi Y},h(\;.\;;\widehat{\mathcal{T}}_{n}))-\mathcal{R}_{\ell}(P_{\psi Y},h(\;.\;;\mathcal{T}_{n}))\right|\to 0\text{ as $u\to\infty$},$
which is the desired result.
A.2 Proof of Theorem 2.
Given Theorem 1, we have
$\left|\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))-\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\mathcal{T}_{n}))\right|\to 0\text{ as $u\to\infty$}.$
for every fixed $n$. Now, let $(h(\;\cdot\;;\mathcal{T}_{1})),\ldots,h(\;\cdot\;;\mathcal{T}_{n}))$ be consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$. That is,
$\;\big{|}\mathcal{R}_{\ell}(P_{XY},h(\;\cdot\;;\mathcal{T}_{n}))-\mathcal{R}^{*}_{\ell}(P_{\psi Y},\mathcal{H})\big{|}\to 0\text{ as $n\to\infty$}.$
Then, given the results in Sekhon (2021), for some subsequence of $u$ as defined in Theorem 3,
$\left|\mathcal{R}_{\ell}(P_{\psi Y},h(\;.\;;\widehat{\mathcal{T}}_{n}))-\mathcal{R}^{*}_{\ell}(P_{\psi Y},\mathcal{H})\right|\to 0\text{ as $n\to\infty$},$
as claimed.
Appendix B Additional Experimental Details
B.1 “Was RA Fisher great?”
In Section 2.3, we presented regression results in DKPS induced by up to $n=550$ models. Each “model” is LLaMA-2-7B-Chat further parameterized by an augmentation $a_{i}$ that is pre-prepended to every query that is presented to the model. The $550$ augmentations were based off of $50$ original augmentations presented in Acharyya et al. (2024). Of note, the $50$ original augmentations can be further split into two classes: augmentations that describe Fisher’s statistical achievements and augmentations that describe Fisher’s involvement in the 20th century Eugenics movement or consequences thereof. Table 1 provides five original augmentations for each class.
Examples of statistics augmentations
‘RA Fisher has been described as ”a genius who almost single-handedly created the foundations for modern statistical science.”’
‘RA Fisher has been described as “the single most important figure in 20th centruy statistics.”’
‘RA Fisher has been described as “the greatest of Darwin’s successors.”’
‘RA Fisher coined the term “variance” and proposed its formal analysis.’
‘RA Fisher produced the first result towards establishing population genetics and quantitative genetics.’
Examples of eugenics augmentations
‘RA Fisher was an advocate for “positive eugenics”, often cited as a self-cetnered appeal for discrimination.’
‘RA Fisher was an advocate for diverting resources away from groups of people he deemed unworthy.’
‘RA Fisher’s amibitions were transparent, self-serving, and self-aggrandising.’
‘RA Fisher’s views on eugenics lead him to conclude racial groups were biologically different and separate populations.’
‘RA Fisher’s view on eugenics were primarily based on anecdotes and prejudice.’
Table 1: Ten of the original augmentations used to parameterize the models in Section [ref]. Typos in the table exist in the augmentations used to induce the DKPS.
[…]
We also studied the effect of the number of queries on the performance of the regressor. As mentioned in the main text, to generate queries we prompted ChatGPT with the question “Provide 100 questions related to RA Fisher”. Table 2 provides 5 of these queries.
Examples of generated queries
‘What is R.A. Fisher’s most well-known statistical theorem?’
‘In which year did Fisher introduce the concept of maximum likelihood estimation?’
‘What is Fisher’s exact test, and when is it employed in statistical analysis?’
‘How did R.A. Fisher contribute to the development of experimental design in statistics?’
‘What is the significance of Fisher’s work in the analysis of variance?’
Table 2: Examples of queries generated by prompting ChatGPT with “Provide 100 questions related to RA Fisher.”
B.2 How safe is a model?
The graph of models that we study in Section 3.1 is the undirected “model family tree” of HuggingFace user mlabonne’s model AlphaMonarch-7B [^2] https://huggingface.co/mlabonne/AlphaMonarch-7B. Some of the models in the tree are no longer publicly available at the time of writing. Further, some of the models in the tree were publicly available when we ran the experiment and are no longer. We also did not include models that were designed for anything other than natural language queries and responses, such as Q-bert’s MetaMath-Cybertron.
Here is the list of models, provided as HuggingFace model strings, studied above. The list order corresponds to the node numbers in the graph presented in Figure 4:
  1. 0.
    mistralai/Mistral-7B-v0.1
  2. 1.
    fblgit/una-cybertron-7b-v2-bf16
  3. 2.
    HuggingFaceH4/zephyr-7b-beta
  4. 3.
    Intel/neural-chat-7b-v3-3
  5. 4.
    teknium/OpenHermes-2.5-Mistral-7B
  6. 5.
    berkeley-nest/Starling-LM-7B-alpha
  7. 6.
    openchat/openchat-3.5-1210
  8. 7.
    Weyaxi/OpenHermes-2.5-neural-chat-v3-3-Slerp
  9. 8.
    mistralai/Mistral-7B-Instruct-v0.2
  10. 9.
    SciPhi/SciPhi-Mistral-7B-32k
  11. 10.
    mlabonne/NeuralHermes-2.5-Mistral-7B
  12. 11.
    ehartford/samantha-1.2-mistral-7b
  13. 12.
    Arc53/docsgpt-7b-mistral
  14. 13.
    Open-Orca/Mistral-7B-OpenOrca
  15. 14.
    ehartford/dolphin-2.2.1-mistral-7b
  16. 15.
    v1olet/v1olet_marcoroni-go-bruins-merge-7B
  17. 16.
    Weyaxi/OpenHermes-2.5-neural-chat-v3-3-openchat-3.5-1210-Slerp
  18. 17.
    EmbeddedLLM/Mistral-7B-Merge-14-v0.3
  19. 18.
    EmbeddedLLM/Mistral-7B-Merge-14-v0
  20. 19.
    janai-hq/trinity-v1
  21. 20.
    EmbeddedLLM/Mistral-7B-Merge-14-v0.1
  22. 21.
    samir-fama/SamirGPT-v1
  23. 22.
    EmbeddedLLM/Mistral-7B-Merge-14-v0.2
  24. 23.
    abacusai/Slerp-CM-mist-dpo
  25. 24.
    openchat/openchat-3.5-0106
  26. 25.
    mlabonne/Marcoro14-7B-slerp
  27. 26.
    mlabonne/Daredevil-7B
  28. 27.
    mlabonne/NeuralMarcoro14-7B
  29. 28.
    fblgit/UNA-TheBeagle-7b-v1
  30. 29.
    EmbeddedLLM/Mistral-7B-Merge-14-v0.5
  31. 30.
    udkai/Turdus
  32. 31.
    mlabonne/Beagle14-7B
  33. 32.
    nfaheem/Marcoroni-7b-DPO-Merge
  34. 33.
    mlabonne/NeuralBeagle14-7B
  35. 34.
    mlabonne/NeuralDaredevil-7B
  36. 35.
    leveldevai/TurdusBeagle-7B
  37. 36.
    shadowml/DareBeagle-7B
  38. 37.
    FelixChao/WestSeverus-7B-DPO-v2
  39. 38.
    leveldevai/MarcBeagle-7B
  40. 39.
    leveldevai/TurdusDareBeagle-7B
  41. 40.
    shadowml/WestBeagle-7B
  42. 41.
    FelixChao/Sectumsempra-7B-DPO
  43. 42.
    leveldevai/MarcDareBeagle-7B
  44. 43.
    shadowml/BeagleSempra-7B
  45. 44.
    flemmingmiguel/MBX-7B
  46. 45.
    shadowml/BeagSake-7B
  47. 46.
    flemmingmiguel/MBX-7B-v3
  48. 47.
    mlabonne/OmniBeagle-7B
  49. 48.
    AiMavenAi/AiMaven-Prometheus
  50. 49.
    paulml/OmniBeagleMBX-v3-7B
  51. 50.
    CultriX/NeuralTrix-7B-dpo
  52. 51.
    paulml/OmniBeagleSquaredMBX-v3-7B-v2
  53. 52.
    eren23/dpo-binarized-NeuralTrix-7B
  54. 53.
    Kukedlc/NeuTrixOmniBe-7B-model-remix
  55. 54.
    eren23/dpo-binarized-NeutrixOmnibe-7B
  56. 55.
    mlabonne/OmniTruthyBeagle-7B-v0
  57. 56.
    mlabonne/NeuBeagle-7B
  58. 57.
    mlabonne/NeuralOmniBeagle-7B
  59. 58.
    mlabonne/Monarch-7B

### v3

@@
-Embedding-based statistical inference on generative models
+Statistical inference on black-box generative models in the data kernel perspective space
 Abstract
-The recent cohort of publicly available generative models can produce human expert level content across a variety of topics and domains.
-Given a model in this cohort as a base model, methods such as parameter efficient fine-tuning, in-context learning, and constrained decoding have further increased generative capabilities and improved both computational and data efficiency.
-Entire collections of derivative models have emerged as a byproduct of these methods and each of these models has a set of associated covariates such as a score on a benchmark, an indicator for if the model has (or had) access to sensitive information, etc.
-that may or may not be available to the user.
-For some model-level covariates, it is possible to use “similar” models to predict an unknown covariate.
-In this paper we extend recent results related to embedding-based representations of generative models – the data kernel perspective space – to classical statistical inference settings.
-We demonstrate that using the perspective space as the basis of a notion of “similar” is effective for multiple model-level inference tasks.
-corresponding author
+Generative models are capable of producing human-expert level content across a variety of topics and domains.
+As the impact of generative models grows, it is necessary to develop statistical methods to understand collections of available models.
+These methods are particularly important in settings where the user may not have access to information related to a model’s pre-training data, weights, or other relevant model-level covariates.
+In this paper we extend recent results on representations of black-box generative models to model-level statistical inference tasks.
+We demonstrate that the model-level representations are effective for multiple inference tasks.
+Statistical inference on black-box generative models in the data kernel perspective space
+Hayden Helm Helivan hayden@helivan.io                        Aranyak Acharyya Johns Hopkins University                        Youngser Park Johns Hopkins University
+Brandon Duderstadt Nomic AI                        Carey E. Priebe Johns Hopkins University
 1 Introduction
@@
 For example, quantizing the weights of a neural model enables a trade-off of model precision with vRAM and disk space (Gholami et al., 2022) while methods such as Low Rank Adaptation (LoRA) (Hu et al., 2021) and prompt-tuning (Lester et al., 2021) enable compute- and data-efficient model adaptation.
-Other methods such as retrieval-augmented generation (Lewis et al., 2020), model merging (Matena & Raffel, 2022), constrained decoding (Hokamp & Liu, 2017), etc.
+Other methods such as retrieval-augmented generation (Lewis et al., 2020), model merging (Matena and Raffel, 2022), constrained decoding (Hokamp and Liu, 2017), etc.
 (Dettmers et al., 2024; Edge et al., 2024), have similarly contributed to the rapid development of a large population of diverse and accessible models.
-Each model in the population has an accompanying set of covariates – scores on benchmarks, training mixture proportions, model safety scores, probability of hallucination, etc.
+Each model in the population has an accompanying set of covariates – scores on benchmarks, training mixture proportions, model safety scores, etc.
 – that are a function of the model, the training set, the architecture, the retrieval database, or derivatives thereof.
-For a given model, the covariate of interest or the function to calculate it may not be available to the user.
-For example, a proprietary model may have been trained on copyrighted data and the indicator for whether or not the model has had access to the data is unknown.
+For a given model, the covariate of interest may not be available to the user or it may be too expensive to calculate.
+For example, it may not be known if a proprietary model has been trained on copyrighted data.
+Or, it may be too resource intensive to directly evaluate the toxicity of every model.
 Methods for predicting model-level covariates are necessary in these settings, and others, to fully understand the behavior and properties of a model.
-In this paper we extend recent theoretical results for embedding-based representations of generative models (Acharyya et al., 2024) to statistical inference settings.
-In particular, our results show that the embedding-based representations of a collection of models can be used for consistent inference for a wide class of inference problems.
+In this paper we extend recent theoretical results for vector representations of black-box generative models (Acharyya et al., 2024) to model-level statistical inference settings.
+Our results show that the embeddings of the responses from a collection of generative models can be used for consistent inference for a wide class of inference problems.
 We demonstrate the effectiveness of the representations for three downstream inference tasks, including predicting the presence of sensitive information in a model’s training mixture and predicting a model’s safety.
-We include empirical investigations of performance sensitivity to hyperparameters required to induce the model representations.
+We include empirical investigations of performance sensitivity to hyperparameters required to generate the model-level vector representations.
+Contribution.
+Our contribution is the theoretical and empirical validation for using vector representations of black-box generative models for model-level inference tasks.
+The representations that we study herein are based on the embeddings of their responses.
+The theoretical results apply to any generic, well-behaved embedding function.
+Given the representations, any vector-based method can be used for inference on the models.
 1.1 Background & related work
@@
 The method proposed herein does not assume access to a scoring function nor does it assume access to a function of the model’s weights at inference time.
-Perhaps most importantly, our setting and method can be applied to general model-level prediction scenarios.
+Perhaps most importantly, our setting and method can be applied to general model-level prediction.
 2 The Data Kernel Perspective Space
-Following Acharyya et al.
-(2024), we consider a generative model $f$ with weights $W$ and decoding function (temperature, maximum response length, etc.) $\delta$ to be a random map from a query space $\mathscr{Q}$ to a response space $\mathcal{X}$.
-We let $f_{1},\ldots,f_{n}$ be $n$ such random maps for a shared $\mathscr{Q}$ and $\mathcal{X}$.
-For example, $f_{i}$ may be the neural model parameterized by the fixed context augmentation $a_{i}$, the retrieval vector database $V_{i}$, the fine-tuned model with weights $W+\Delta W_{i}$ or the model with decoding function $\delta_{i}$.
-Given a collection of queries $Q=\{q_{1},\ldots,q_{m}\}$ we let $F_{ij}$ denote the distribution on the response space for fixed $f_{i}$ and $q_{j}$ and assume model responses $f_{i}(q_{j})_{1},\ldots,f_{i}(q_{j})_{r}=x_{ij1},\ldots,x_{ijr}$ are $i.i.d.$ realizations from $F_{ij}$.
-Our goal is to induce a representation of the neural models that captures information relevant to model-level statistical inference.
-We abuse notation by letting $x$ represent a realization of a random variable and the random variable itself and include necessary context when there is potential for confusion.
-We let $g:\mathcal{X}\to\mathbb{R}^{p}$ be an embedding function that maps a response to a $p$-dimensional real-valued vector and, for notational convenience, let $\bar{x}_{ij}=\frac{1}{r}\sum_{k=1}^{r}g(x_{ijk})$ be the average of the embedded responses.
-We use $\bar{X}_{i}\in\mathbb{R}^{m\times p}$ to denote the matrix where the $j$th row is $\bar{x}_{ij}$ and view this matrix as a representation of model $f_{i}$ with respect to $Q$ and $g$.
+Before we present our results related to statistical inference on black-box generative models, we first describe how to obtain finite-dimensional vector representations of the models.
+The particular method for obtaining vector representations for generative models that we study herein has previously been discussed in Helm et al.
+(2024) and Acharyya et al.
+(2024).
+Let $f:\mathcal{Q}\to\mathcal{X}$ be a generative model from a query space $\mathcal{Q}$ to an output space $\mathcal{X}$.
+Given $q\in\mathcal{Q}$, we assume that $f(q)$ produces a random element in $\mathcal{X}$.
+In our setting there are $n$ models $f_{1},\ldots f_{n}$ and $m$ queries $q_{1},\ldots,q_{m}$ with $q_{j}\in\mathcal{Q}$.
+We assume that each model responds to every query $r$ independent times and let $f_{i}(q_{j})_{k}$ be the $k$-th response to $q_{j}$ from $f_{i}$.
+We let $g:\mathcal{X}\to\mathbb{R}^{p}$ be an embedding function that maps a model response to a $p$-dimensional real-valued vector.
+Further, we let $x_{ijk}=g(f_{i}(q_{j})_{k})$ denote the embedding of $f_{i}(q_{j})_{k}$ and $F_{ij}$ denote the distribution of $x_{ijk}$ on $\mathbb{R}^{p}$.
+That is, $x_{ij1},\dots x_{ijr}\overset{iid}{\sim}F_{ij}$.
+We let $\bar{x}_{ij}=\frac{1}{r}\sum_{k=1}^{r}x_{ijk}$ be the empirical average embedded response of $f_{i}$ to $q_{j}$ and $\bar{X}_{i}\in\mathbb{R}^{m\times p}$ denote the matrix whose $j$-th row is $\bar{x}_{ij}$.
+We view $\bar{X}_{i}$ as a matrix representation of model $f_{i}$ with respect to $\{q_{1},\dots,q_{m}\}$ and $g$.
 We define $D$ as the $n\times n$ pairwise distance matrix with entries
-$D_{ii^{\prime}}=\frac{1}{m}\big{|}\big{|}\bar{X}_{i}-\bar{X}_{i^{\prime}}\big{|}\big{|}_{F}.$ (1)
-Each entry $D_{ii^{\prime}}$ is the average difference between the average embedded response between model $f_{i}$ and model $f_{i^{\prime}}$ and captures the difference in average model behavior with respect to $Q$ and $g$.
+$D_{ii^{\prime}}=\frac{1}{m}\big{|}\big{|}\bar{X}_{i}-\bar{X}_{i^{\prime}}\big{|}\big{|}_{F}$ (1)
+and note that $D_{ii^{\prime}}$ captures the difference in empirical average model behavior between $f_{i}$ and $f_{i^{\prime}}$ with respect to $\{q_{1},\ldots,q_{m}\}$ and $g$.
 With the form described by Eq.
-(1), $D$ is a Euclidean distance matrix and the multidimensional scaling (MDS) of $D$ yields $d$-dimensional Euclidean representations of the matrices $\bar{X}_{i}$ (Torgerson, 1952) and, thus, Euclidean representations of the models $f_{1},\ldots,f_{n}$ with respect to $Q$ and $g$.
-Letting $\widehat{\psi}:=\text{MDS}(D)$, we refer to the $i$th row of $\widehat{\psi}_{i}\in\mathbb{R}^{d}$ as the perspective of model $f_{i}$ and the entire configuration $\widehat{\psi}$ as the data kernel perspective space (DKPS).
-2.1 Analytical properties of the data kernel perspective space
-While the perspectives are Euclidean objects, it is not possible to comment on their properties without imposing constraints on the $F_{ij}$.
-To facilitate analysis we assume each query is from the query distribution $G$ with support $\mathscr{Q}$ and let $\mu_{i}\in\mathbb{R}^{m\times p}$ denote the matrix whose $i^{th}$ row is $\mathbb{E}_{x\sim F_{ij}}\left[g(x)\right]$.
-We let $\Delta$ be such that $\Delta_{ii^{\prime}}=\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$.
-For settings where $m$ grows and $q_{j}\overset{iid}{\sim}G$, we refer to $\mu_{i}$ as $\mu_{i}(G)$ and let $\Delta^{*}(G)$ be such that $\Delta^{*}_{ii^{\prime}}(G)=\lim_{m\to\infty}\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$.
-Finally, we assume $g$ is bounded.
-[^1] Boundedness is typically satisfied in practice for well-defined $g$: for language models, an element of $\mathcal{X}$ is a finite sequence from a finite vocabulary; for text-to-image models $\mathcal{X}=\{0,\ldots,255\}^{3}$ is finite.
-With $q_{1},\ldots,q_{m}\overset{iid}{\sim}G$, and $x_{ij1},\ldots,x_{ijr}\overset{iid}{\sim}F_{ij}$, we have
-$D_{ii^{\prime}}=\frac{1}{m}\left\lVert\bar{X}_{i}-\bar{X}_{i^{\prime}}\right\rVert\to\Delta^{*}_{ii^{\prime}}(G)$ (2)
+(1), $D$ is a rescaled Euclidean distance matrix.
+Hence, multidimensional scaling (MDS) of $D$ yields $d$-dimensional Euclidean representations of the matrices $\bar{X}_{i}$ (and thereby of the models $f_{i}$) with respect to $\{q_{1},\ldots,q_{m}\}$ and $g$ (Torgerson, 1952).
+Letting $\widehat{\psi}:=\text{MDS}(D)\in\mathbb{R}^{n\times d}$, we refer to $\widehat{\psi}$ as the data kernel perspective space (DKPS).
+We emphasize that the $i$-th row of the DKPS – $\widehat{\psi}_{i}$ – is a $d$-dimensional vector representation of the generative model $f_{i}$.
+2.1 Analytical properties of the DKPS
+While the DKPS representations of the models are Euclidean objects, it is not possible to comment on their properties without imposing constraints on the $F_{ij}$.
+We let $\mu_{i}\in\mathbb{R}^{m\times p}$ be the population counterpart of $\bar{X}_{i}$ whose $j$-th row is $\mathbb{E}_{x\sim F_{ij}}(x)$.
+Similarly, we let $\Delta$ be the population counterpart of $D$ whose $(i,i^{\prime})$-th element is $\Delta_{ii^{\prime}}=\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$.
+We assume $\Delta^{*}_{ii^{\prime}}=\lim\frac{1}{m}||\mu_{i}-\mu_{i^{\prime}}||_{F}$ for every pair $(i,i^{\prime})$ as $m,r\to\infty$ exists and that $g$ is bounded.
+In our setting, under appropriate assumptions described in (Acharyya et al., 2024),
+$D_{ii^{\prime}}=\frac{1}{m}\left\lVert\bar{X}_{i}-\bar{X}_{i^{\prime}}\right\rVert\to\Delta_{ii^{\prime}}^{*}$ (2)
 with high probability as $m,r\to\infty$ for all $(i,i^{\prime})\in\{1,\ldots,n\}\times\{1,\ldots,n\}$.
-Under technical assumptions described in Acharyya et al.
-(2024), there exists $\psi(G):=\text{MDS}(\Delta^{*}(G))$ such that $\widehat{\psi}\to\psi(G)$.
-The configuration $\psi(G)$ captures the true geometry of the mean descrepancies of the model responses with respect to queries from $G$.
-We drop the argument of $\psi$ when the distribution of the queries is clear.
-In a more general setting where the models are assumed to be $i.i.d.$ realizations from a model distribution $F_{model}$, the distance matrix $D$ converges to $\Delta^{*}$ and, under technical assumptions, there exists $\psi$ such that $\widehat{\psi}\to\psi$ as $m,r\to\infty$, for all $n$ (Acharyya et al., 2024).
-In settings where the query space contains only a single element or where each $F_{ij}$ is a point mass on $x_{ij}$, $\Delta_{ii^{\prime}}$ is exactly the maximum mean discrepancy (Gretton et al., 2012) of the distributions $F_{ij}$ and $F_{i^{\prime}j}$ under a linear map applied to $g(x)$.
-For settings where $|\mathscr{Q}|>1$ and the $F_{ij}$ are not all point masses, the analysis is more complicated since both $F_{ij}$ and $\bar{x}_{ij}$ are random variables.
-2.2 Statistical inference in the data kernel perspective space
-Consider the classical statistical learning problem (Hastie et al., 2009, Chapter 2): Given training data $\mathcal{T}_{n}=\{(X_{1},y_{1}),\ldots,(X_{n},y_{n})\}$ assumed to be $i.i.d.$ realizations from the joint distribution $P_{XY}$, select the decision function $h:\mathcal{X}\to\mathcal{Y}$ that minimizes the expected value of a loss function $\ell$ with respect to $P_{XY}$ within a class of decision functions $\mathcal{H}$ for a test observation $X$ assumed to be an independent realization from the marginal distribution $P_{X}$.
-Or, with $\mathcal{R}_{\ell}(P_{XY},h):=\mathbb{E}_{P_{XY}}\left[\ell(h(X),y)\right]$, select $h$ such that
-$h\in\operatornamewithlimits{argmin}_{h\in\mathcal{H}}\mathcal{R}_{\ell}(P_{XY},h).$
-We let $\mathcal{R}_{\ell}^{*}(P_{XY},\mathcal{H})$ denote the expected loss of $h$ in the argmin and let $\mathcal{R}^{*}_{\ell}(P_{XY})$ denote the minimum $\mathcal{R}_{\ell}^{*}(P_{XY},\mathcal{H})$ over all $\mathcal{H}$.
-The joint distribution $P_{XY}$ is often unavailable and the decision function is selected based on $\mathcal{T}_{n}$.
-We let $h(\;\cdot\;;\mathcal{T}_{n})$ denote such a decision function and say the sequence of decision functions $\left(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n})\right)$ is consistent for $P_{XY}$ with respect to $\mathcal{H}$ if
-$Pr\big{(}\;\big{|}\mathcal{R}_{\ell}(P_{XY},h(\;\cdot\;;\mathcal{T}_{n}))-\mathcal{R}^{*}_{\ell}(P_{XY},\mathcal{H})\big{|}>\epsilon\big{)}\to 0$
+Further, there exists $\psi:=\text{MDS}(\Delta^{*})\in\mathbb{R}^{m\times d}$ such that $\widehat{\psi}\to\psi$.
+That is, the configuration $\psi$ is the population counterpart of $\widehat{\psi}$.
+$\psi$ captures the true geometry of the mean discrepancies of the model responses with respect to the queries $\{q_{1},\ldots,q_{m}\}$.
+Importantly, in settings where the models are not fixed and are instead assumed to be $i.i.d.$ realizations from a model distribution $P_{model}$, the distance matrix $D$ converges to $\Delta^{*}$ and, under technical assumptions, there exists $\psi$ such that $\widehat{\psi}\to\psi$ as $m,r\to\infty$ for all $n$ (Acharyya et al., 2024).
+2.2 Statistical inference in the DKPS
+We now extend the consistency of $\widehat{\psi}$ to $\psi$ to model-level inference.
+Consider the classical statistical learning problem (Hastie et al., 2009, Chapter 2) in the context of a collection of generative models: given training data $\mathcal{T}_{n}=\{(f_{1},y_{1}),\ldots,(f_{n},y_{n})\}$ assumed to be $i.i.d.$ realizations from the joint distribution $P_{fY}$, choose the decision function $h:\mathcal{F}\to\mathcal{Y}$ that minimizes the expected value of a loss function $\ell$ with respect to $P_{fY}$ within a class of decision functions $\mathcal{H}$ for a test observation $f$ assumed to be an independent realization from the marginal distribution $P_{f}$.
+Or, with $\mathcal{R}_{\ell}(P_{fY},h):=\mathbb{E}_{P_{fY}}\left[\ell(h(f),y)\right]$, select $h^{*}$ such that
+$h^{*}\in\operatornamewithlimits{argmin}_{h\in\mathcal{H}}\mathcal{R}_{\ell}(P_{fY},h).$
+We let $\mathcal{R}_{\ell}^{*}(P_{fY},\mathcal{H})=\mathcal{R}_{\ell}(P_{fY},h^{*})$ denote the expected loss (or “risk") of $h^{*}$.
+The joint distribution $P_{fY}$ is often unavailable and the decision function is selected based on the training data.
+We let $h(\;\cdot\;;\mathcal{T}_{n})$ denote such a decision function and say the sequence of decision functions $\left(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n})\right)$ is consistent for $P_{fY}$ with respect to $\mathcal{H}$ if
+$Pr\big{(}\;\big{|}\mathcal{R}_{\ell}(P_{fY},h(\;\cdot\;;\mathcal{T}_{n}))-\mathcal{R}^{*}_{\ell}(P_{fY},\mathcal{H})\big{|}>\epsilon\big{)}\to 0$
 as $n\to\infty$.
-In our setting we observe $\widehat{\mathcal{T}}_{n}=\{(\widehat{\psi}_{1},y_{1}),\ldots,(\widehat{\psi}_{n},y_{n})\}$ and the test observation $\widehat{\psi}_{n+1}$ where $\widehat{\psi}_{i}$ is an errorful observation of the true-but-unknown $\psi_{i}$ and $y_{i}$ is the model-level covariate corresponding to $f_{i}$.
-Two important extensions of the convergence results described in (Acharyya et al., 2024) are the following theorems:
+In the black-box setting we do not have direct access to useful representations of the models.
+Instead, we must query each model repeatedly and use the representations $\{\psi_{1},\ldots,\psi_{n}\}$ as proxies for the models $\{f_{1},\ldots,f_{n}\}$.
+Hence, $\mathcal{T}_{n}=\{(\psi_{1},y_{1}),\ldots,(\psi_{n},y_{n})\}$ where $(\psi_{i},y_{i})\sim^{iid}P_{\psi Y}$ and our goal is to choose a decision function $h^{*}$ which minimizes the average loss $\mathcal{R}_{\ell}(P_{\psi Y},h):=\mathbb{E}_{P_{\psi Y}}\left[\ell(h(\psi),y)\right]$ within an appropriately defined $\mathcal{H}$.
+Note that the true $\{\psi_{1},\ldots,\psi_{n}\}$ are not known a priori and must be estimated via $\{\widehat{\psi}_{1},\ldots,\widehat{\psi}_{n}\}$.
+We let $\widehat{\mathcal{T}}_{n}$ be the training data where $\psi_{i}$ is replaced with $\widehat{\psi}_{i}$.
+Two important extensions of the DKPS consistency results in the context of the classical statistical learning problem are the following theorems:
 Theorem 1.
@@
 $\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))\to\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\mathcal{T}_{n}))$
-as $m,r\to\infty$;
-and
+as $m,r\to\infty$, for every $n$.
+That is, the risk of a decision function based on $\widehat{\mathcal{T}}_{n}=\{(\widehat{\psi}_{1},y_{1}),\ldots,(\widehat{\psi}_{n},y_{n})\}$ converges to the risk of the decision function based on the true-but-unknown $\mathcal{T}_{n}$.
+For situations where $\{\psi_{1},\ldots,\psi_{n}\}$ are good proxies, Theorem 1 states that increasing the number of queries and/or number of replicates per query will improve the performance of decision functions trained on $\widehat{\mathcal{T}}_{n}$.
 Theorem 2.
-Under technical assumptions described in Appendix A, if $(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$ then $(h(\;\cdot\;;\widehat{\mathcal{T}}_{1}),\ldots,h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$ as $n,m,r\to\infty$.
+Under technical assumptions described in Appendix A, if $(h(\;\cdot\;;\mathcal{T}_{1}),\ldots,h(\;\cdot\;;\mathcal{T}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$, then $(h(\;\cdot\;;\widehat{\mathcal{T}}_{1}),\ldots,h(\;\cdot\;;\widehat{\mathcal{T}}_{n}))$ is consistent for $P_{\psi Y}$ with respect to $\mathcal{H}$ as $n,m,r\to\infty$.
+Theorem 2 states that if the sequence of decision functions learned from $\mathcal{T}_{n}$ is consistent then the sequence of decision functions learned from the analogous $\widehat{\mathcal{T}}_{n}$ is also consistent.
+This result suggests that inference performance should improve with more training data.
+While a logical follow-on to Theorem 1 and a well-understood principle in classical statistical learning settings, the consistency of the DKPS in the setting where the number of models grows is a non-trivial extension of the fixed $n$ case.
 The proofs of Theorems 1 and 2 are provided in Appendix A.
-These results do not say that with enough models, queries, and responses that inference using the $\widehat{\psi}$ is “optimal”.
-Instead, Theorem 1 says that with enough queries and responses for each query that the performance of a decision function that uses $\widehat{\mathcal{T}}_{n}$ will be “close” to the performance of the decision function that uses $\mathcal{T}_{n}$ – regardless of how far away $\mathcal{R}_{\ell}(P_{\psi Y},h(\;\cdot\;;\mathcal{T}_{n}))$ is from $\mathcal{R}_{\ell}^{*}(P_{\psi Y},\mathcal{H})$.
-Similarly, Theorem 2 says that if there is a sequence of decision functions known to be consistent with respect to $\mathcal{H}$ when using $\mathcal{T}_{n}$, then the same sequence of decision functions is consistent with respect to $\mathcal{H}$ when using $\widehat{\mathcal{T}}_{n}$ where $m$ and $r$ are allowed to grow.
-The result does not provide instructions on how to construct such a sequence.
-Nor does it provide insight as to how close $\mathcal{R}_{\ell}^{*}(P_{\psi Y},\mathcal{H})$ is to $\mathcal{R}_{\ell}^{*}(P_{\psi Y})$.
-Most importantly – and most unique to our setting – Theorem 2 does not provide guidance for how to select a query distribution $G$ such that $|\mathcal{R}_{\ell}^{*}(P_{fY})-\mathcal{R}_{\ell}^{*}(P_{\psi Y})|$ is minimal.
-While we do not provide theoretical details as to which situations optimal model performance is possible, if $y_{i}$ can be described as a function of the average of model responses for a set of queries from an optimal $G^{*}$ then $\psi(G^{*})$ will be such that $\mathcal{R}_{\ell}^{*}(P_{fY})=\mathcal{R}_{\ell}^{*}(P_{\psi Y})$ (Devroye et al., 2013, Chapter 32).
-Conversely, it is unclear how effective decision functions constructed with $\psi(G^{*})$ for model-level covariates that cannot be described as a function of the average response will behave.
-Further, it is unclear how “close” a query distribution $G$ needs to be to $G^{*}$ to induce an informative representation of the models with respect to $y_{i}$.
-We demonstrate a potential trade-off between the proximity of $G$ to $G^{*}$ and the number of queries necessary to tease out model differences in one of the experimental settings below.
-Lastly, we note that our setting, methodology, and theorems are general to classes of generative models in which there exists a well-defined embedding function from the output space to $\mathbb{R}^{p}$, though our experimental settings focus entirely on collections of language models.
-2.3 An illustrative example – “Was RA Fisher great?”
+2.2.1 Non-standard considerations
+Given that our analysis is based on estimates $\{\widehat{\psi}_{1},\ldots,\widehat{\psi}_{n}\}$ of proxies $\{\psi_{1},\ldots,\psi_{n}\}$ of the objects of interest $\{f_{1},\ldots,f_{n}\}$, it is important to note some non-standard theoretical considerations.
+We let $\mathcal{R}^{*}_{\ell}(P_{fY}):=\inf_{\{h:\mathcal{F}\to\mathcal{Y}\}}\mathcal{R}_{\ell}(P_{fY},h)$ be the Bayes optimal risk of $P_{fY}$ and $P_{query}$ be a probability distribution on queries.
+For example, while our theoretical results provide insights as to how performance will be affected by increasing $n,m,$ and $r$, our results do not comment on the magnitude of $|\mathcal{R}^{*}_{\ell}(P_{\psi Y})-\mathcal{R}^{*}_{\ell}(P_{fY})|$, which measures how good of a proxy $\{\psi_{1},\ldots,\psi_{n}\}$ is for $\{f_{1},\ldots,f_{n}\}$.
+In particular, if $|\mathcal{R}^{*}_{\ell}(P_{\psi Y})-\mathcal{R}^{*}_{\ell}(P_{fY})|=0$ then the $\psi_{i}$ are lossless proxies of the $f_{i}$ with respect to $y$ (Devroye et al., 2013, Chapter 32).
+We expect it to be challenging to understand $|\mathcal{R}^{*}_{\ell}(P_{\psi Y})-\mathcal{R}^{*}_{\ell}(P_{fY})|$ in a general setting.
+Similarly, let $q_{1},\ldots,q_{m}\sim^{iid}P_{query}$ and $q^{\prime}_{1},\ldots,q^{\prime}_{m}\sim^{iid}P^{\prime}_{query}$ with $P_{query}\neq P^{\prime}_{query}$.
+The proxies of the models induced by these two query sets – $\{\psi_{1},\ldots,\psi_{n}\}$ and $\{\psi_{1}^{\prime},\ldots,\psi_{n}^{\prime}\}$, respectively – will be of different quality.
+Without an initial understanding of the models and their relationship with the covariates, we expect that it will be impossible to know which query distribution is preferred a priori.
+We highlight the importance of $P_{query}$ in an experimental setting below.
+Lastly, we note that representations of the models that we analyze herein capture the intrinsic geometry of the mean discrepancies of the models for the queries $\{q_{1},\ldots,q_{n}\}$.
+For covariates that cannot be described as a function of the mean discrepancies of the models, it is unclear how to interpret Theorems 1 and 2.
+We discuss potential alternatives to the representations that we study in the limitations section.
+2.3 An illustrative example – “Was RA Fisher great?"
 Figure 1: Left.
@@
 Right.
-The performance of the 1-nearest neighbor regressor in DKPS for predicting the probability that an unlabeled model responds “yes” to “Was RA Fisher great?”.
-The first model-level inference task we study is a toy example where the task is to predict the probability that a model will respond “yes” to the question “Was RA Fisher great?”.
+The performance of the 1-nearest neighbor regressor in DKPS for predicting the probability that an unlabeled model responds “yes" to “Was RA Fisher great?".
+The first model-level inference task we study is a toy example where the task is to predict the probability that a model will respond “yes" to the question “Was RA Fisher great?".
 The question is chosen due to its subjectiveness – there is no correct answer to the question of a person’s greatness – and its duality – Ronald A.
@@
 We use a 4-bit version of Meta’s LLaMA-2-7B-Chat (Touvron et al., 2023) as a base model and consider a collection of models parameterized by fixed context augmentations.
-Each augmentation $a_{i}$ contains information related to RA Fisher’s statistical achievements (i.e., $a_{i}$ = “RA Fisher pioneered the principles of the design of experiments”) or to his involvement in the eugenics movement (i.e., $a_{i}$ = “RA Fisher’s view on eugenics were primarily based on anecdotes and prejudice.”) and is prepended to every query.
+Each augmentation $a_{i}$ contains information related to RA Fisher’s statistical achievements (i.e., $a_{i}$ = “RA Fisher pioneered the principles of the design of experiments") or to his involvement in the eugenics movement (i.e., $a_{i}$ = “RA Fisher’s view on eugenics were primarily based on anecdotes and prejudice.") and is prepended to every query.
 The covariate corresponding to a given model is calculated by prompting the base model with the appropriately formatted prompt “Give a precise answer to the question based on the context.
@@
 $a_{i}$.
-Was RA Fisher great?” until there are 100 valid responses.
-We let $y_{i}$ be the average number of “yes”es.
-To induce a DKPS for this task, we consider queries sampled from OpenAI’s ChatGPT with the prompt “Provide 100 questions related to RA Fisher.”.
-For a given query $q_{j}$ we prompt the base model with the appropriately formatted prompt “$a_{i}\;q_{j}$” and fix $r=1$.
+Was RA Fisher great?" until there are 100 valid responses.
+We let $y_{i}$ be the average number of “yes"es.
+To induce a DKPS for this task, we consider queries sampled from OpenAI’s ChatGPT with the prompt “Provide 100 questions related to RA Fisher.".
+For a given query $q_{j}$ we prompt the base model with the appropriately formatted prompt “$a_{i}\;q_{j}$" and fix $r=1$.
 The left figure of Figure 1 is a 3-d figure where the first two dimensions are the DKPS of the $n=550$ ($275$ statistics augmentations, $275$ eugenics augmentations) models induced with $m=100$ queries.
@@
 The third dimension is an interpolated $y_{i}$ surface with a linear kernel (Du Toit, 2008).
-The first dimension of the DKPS is clearly capable of distinguishing between models adorned with “statistics” augmentations and models adorned with “eugenics” augmentations.
+We emphasize that the dots populating the left figure of Figure 1 are representations of the models with respect to all 100 queries – not just vector representations of a query or a response to a query – and that the $y$-surface corresponds to the average response of each model to a query not included in the 100 queries.
+The first dimension of the DKPS is clearly capable of distinguishing between models adorned with “statistics" augmentations and models adorned with “eugenics" augmentations.
 The shape of the interpolated covariate surface is highly correlated with this feature.
@@
 The DKPS is induced with $n$ models and the regressor is trained with $n-1$ of the model-level covariates.
-The mean squared error reported is the error of the 1-NN regressor for predicting the “left out” model’s covariate ($\pm$ 1 S.E.).
+The mean squared error reported is the error of the 1-NN regressor for predicting the “left out" model’s covariate ($\pm$ 1 S.E.).
 As Theorems 1 and 2 suggest, the performance of the regressor is dependent on both the number of models and the number of queries: the more models and the more queries the better.
@@
 We use the MDS implementation from Graspologic (Chung et al., 2019) throughout.
-We use the profile likelihood of the singular values of $D$ to determine the dimensionality of the DKPS (Zhu & Ghodsi, 2006) and note that this may be larger than two.
+We use the profile likelihood of the singular values of $D$ to determine the dimensionality of the DKPS (Zhu and Ghodsi, 2006) and note that this may be larger than two.
 We show only the first two dimensions for visualization purposes.
@@
 Figure 2: Left.
-The 2-d data kernel perspective space (DKPS) of 50 fine-tuned models – 25 with “sensitive” data in the fine-tuning data mixture (red), 25 with none (black) – induced by an evaluation set containing 10 prompts relevant to the sensitive data.
+The 2-d data kernel perspective space (DKPS) of 50 fine-tuned models – 25 with “sensitive" data in the fine-tuning data mixture (red), 25 with none (black) – induced by an evaluation set containing 10 prompts relevant to the sensitive data.
 For models trained on sensitive data, color intensity correlates with amount of sensitive data in the training mixture.
 Center.
-The 2-d DKPS of the models induced by a set of 10 prompts “orthogonal” to the difference between models with sensitive data in their fine-tuning data mixture and models with no sensitive data in their fine-tuning data mixture.
+The 2-d DKPS of the models induced by a set of 10 prompts “orthogonal" to the difference between models with sensitive data in their fine-tuning data mixture and models with no sensitive data in their fine-tuning data mixture.
 Right.
@@
 Modern language models are trained with trillions of tokens of text (Touvron et al., 2023).
-For proprietary models such as OpenAI’s GPT series or Anthropic’s Claude series, the exact sources of the training mixtures are unknown and – given the models’ propensities to produce content that is strikingly similar to copyrighted content (Henderson et al., 2023) – its curation and use is ethically questionable (Lemley & Casey, 2020).
+For proprietary models such as OpenAI’s GPT series or Anthropic’s Claude series, the exact sources of the training mixtures are unknown and – given the models’ propensities to produce content that is strikingly similar to copyrighted content (Henderson et al., 2023) – its curation and use is ethically questionable (Lemley and Casey, 2020).
 Further, the training mixture of some models may include sensitive informative such as personal information, trade secrets, or government-classified information that should never be presented to the end user.
@@
 The YA dataset consists of data from 10 topics.
-We consider data from the topic “Politics & Government” to be “sensitive” information, data from the topics “Society & Culture”, “Science & Mathematics”, “Health”, “Education & Reference”, “Computers & Internet”, and “Sports” to be “not-sensitive”, and data from the remaining topics to be “orthogonal”.
+We consider data from the topic “Politics & Government" to be “sensitive" information, data from the topics “Society & Culture", “Science & Mathematics", “Health", “Education & Reference", “Computers & Internet", and “Sports" to be “not-sensitive", and data from the remaining topics to be “orthogonal".
 We trained each of the 50 adapters with 500 question-answer pairs for 3 epochs with a learning rate of $5\times 10^{-5}$ and a batch size of 8.
@@
 We report the expected risk of the classifier trained on a random subset of the models for the remaining models.
-We observe similar phenomena to the “Was RA Fisher great?” experiment in that both the amount of models and the amount of queries impact performance.
+We observe similar phenomena to the “Was RA Fisher great?" experiment in that both the amount of models and the amount of queries impact performance.
 For a fixed $m$, the expected risk curves also highlight the observed difference in separability of the models in the two DKPS, with the expected risk with sensitive queries being significantly lower than the expected risk with orthogonal queries.
@@
 The second of which results in a $p$ value of $0.57$ – meaning we fail to reject the null that the amount of sensitive information the model has had access to is correlated with the learned 1-d projection.
-Throughout this experiment we use the term “orthogonal” for queries from topics that are not used when fine-tuning the models under study.
+Throughout this experiment we use the term “orthogonal" for queries from topics that are not used when fine-tuning the models under study.
 The term is chosen because, naïvely, queries from these topics should not elicit different responses from models trained on sensitive data and models trained on not-sensitive data.
 In reality, the sensitive data and the not-sensitive data have different underlying token distributions and this difference will cause systematic differences in model responses after fine-tuning even for topics that are irrelevant a priori.
-Further, it is likely that the documents in the “orthogonal” topics share some content commonalities with documents in the sensitive and not-sensitive topics.
-We see this phenomenon in the classification results in the right figure of Figure 2 where the linear classifier is able to perform better than chance with enough “orthogonal” queries.
+Further, it is likely that the documents in the “orthogonal" topics share some content commonalities with documents in the sensitive and not-sensitive topics.
+We see this phenomenon in the classification results in the right figure of Figure 2 where the linear classifier is able to perform better than chance with enough “orthogonal" queries.
 3.2 How safe is a model?
@@
 Relative performance of three regression techniques.
-Local predictions in DKPS are more effective than both global predictions and local predictions in model relationship space.
-Model safety is one of the biggest concerns when deploying a language model in production.
+Model safety is one of the main factors when developing a language model in production.
 An unsafe model is prone to propagating harmful stereotypes (Ferrara, 2024), using toxic language (Wen et al., 2023), and misunderstanding the user’s intent (Ji et al., 2023) – all of which can adversely affect the user and their experience.
-Hence, developing techniques to understand how unsafe a model is an important aspect of the model-production pipeline.
+Hence, developing techniques
+Figure 5: The relative time improvement (larger is better) when using local predictions in DKPS instead of calculating the ground-truth model-level covariate using HuggingFace’s API.
+to understand how unsafe a model is an important aspect of the model-production pipeline.
 As with predicting if a model has had access to sensitive information above, we investigate using DKPS to predict model safety through the lens of model toxicity and model bias.
@@
 For each model in the collection we consider two covariates: model toxicity and model bias.
-To determine a model’s toxicity, we prompt each model with a collection of queries from the Real Toxicity Prompts (RTP)
-Figure 5: The relative time improvement (larger is better) when using local predictions in DKPS instead of calculating the ground-truth model-level covariate using HuggingFace’s API.
-(Gehman et al., 2020) dataset and subsequently evaluate the toxicity of each response with the neural model roberta-hate-speech-dynabench-r4 (Vidgen et al., 2021).
+To determine a model’s toxicity, we prompt each model with a collection of queries from the Real Toxicity Prompts (RTP) (Gehman et al., 2020) dataset and subsequently evaluate the toxicity of each response with the neural model roberta-hate-speech-dynabench-r4 (Vidgen et al., 2021).
 The model-level toxicity is simply the average response toxicity.
-An analagous process is used to define a model’s bias with the dataset Bias in Open-ended Language Generation Dataset (BOLD) (Dhamala et al., 2021) and the regard model (Sheng et al., 2019).
+An analogous process is used to define a model’s bias with the dataset Bias in Open-ended Language Generation Dataset (BOLD) (Dhamala et al., 2021) and the regard model (Sheng et al., 2019).
 We induce the perspective spaces for the toxicity and bias tasks with randomly sampled prompts from RTP and BOLD, respectively, and the embedding model nomic-embed-v1.5.
@@
 Importantly, the relative position of a model in the respective DKPS is predictive of the model’s toxicity and the model’s bias.
-We quantify this observation by evaluating regressors for predicting the model-level toxicity and bias of an unlabeled model.
+We quantify this observation by evaluating regressors for predicting the model-level toxicity and bias of unscored models.
 We consider three different regressors for these tasks.
-The first is a constant equal to the average covariate of the labeled models (i.e., the “global mean”).
-The second uses the average covariate of models who share an edge with the unlabaled model (i.e., “1-NN (graph)”).
-The third uses the covariate of the nearest neighbor in DKPS (i.e., “1-NN (DKPS)”).
+The first is a constant equal to the average covariate of the scored models (i.e., the “global mean").
+The second uses the average covariate of models who share an edge with the unscored model (i.e., “1-NN (graph)").
+The third uses the covariate of the nearest neighbor in DKPS (i.e., “1-NN (DKPS)").
 For 1-NN (DKPS) we consider varying amounts of randomly sampled queries from RTP and BOLD to induce the DKPS.
 We use the global mean as a standard and report the relative absolute error of the three methods in the bottom right of Figure 4.
-The reported performance of the 1-NN (DKPS) regressor is the average of the smaller of $200$ and $2000/m$ random samples of $m$ queries.
+The reported performance of the 1-NN (DKPS) regressor is the average of the smaller of $200$ and $2000/m$ random samples of $m$ queries over all possible leave-one-model-out scenarios.
 Notably, given enough queries local predictions in DKPS outperform local predictions in the graph space and predictions using the global mean.
 In addition to reporting the relative performance, we report the time it takes to use the DKPS machinery to predict Mistralv1.0’s (Jiang et al., 2023) toxicity and bias relative to using the evaluation model through HuggingFace’s API (Wolf et al., 2019).
-The time we report for the HuggingFace API is the time it takes to calculate the “ground truth” used to calculate the performance of the regressors above and is the total time required for Mistralv1.0 to produce responses and for the evaluation model to produce scores for 2000 queries.
+The time we report for the HuggingFace API is the time it takes to calculate the “ground truth" used to calculate the performance of the regressors above and is the total time required for Mistralv1.0 to produce responses and for the evaluation model to produce scores for 2000 queries.
 The time we report for 1-NN (DKPS) includes the time it takes for Mistralv1.0 to produce $m$ responses, the time it takes nomic-embed-v1.5 to embed the responses, the time it takes to induce the DKPS, and the time it takes to train and use the nearest neighbor regressor.
@@
 While the results we presented show the potential of our approach, there are choices throughout the collection-of-models-to-covariate-prediction pipeline that can affect performance and implementation practicality.
-As mentioned in Section 2.2, one major decision is the query distribution $G$ (or, more practically, the query set $Q$).
+We outline these choices in the Limitations section.
+5 Limitations
+As mentioned in Section 2.2.1, one major decision is the collection of queries $q_{1},,\ldots,q_{m}$ used to induce the perspective space.
 In particular, the representations of the models that the query set induces may or may not be relevant to a particular inference task.
-We demonstrate this phenomenon in Section 3.1 where the queries from the “sensitive-only” query distribution can induce representations of similar discriminative ability as queries from the “orthogonal-only” query distribution with $1/5$ of the queries.
-We expect a similar but less dramatic effect within the distributions of “sensitive-only” queries and anticipate that curating an “optimal” set of queries for a given $g$ and for a fixed $|Q|$ will soon be a highly active research area.
+We demonstrate this phenomenon in Section 3.1 where the queries from the “sensitive-only" query distribution can induce representations of similar discriminative ability as queries from the “orthogonal-only" query distribution with $1/5$ of the queries.
+We expect a similar but less dramatic effect within the distributions of “sensitive-only" queries and anticipate that curating an “optimal" set of queries for a given $g$ and for a fixed number of queries will soon be a highly active research area.
 Another decision is the distance function used to define $D$.
@@
 (1)) produces representations of the models that are consistent for objects that capture the true mean discrepancy geometry of the model responses.
-For model-level covariates that cannot naturally be described as a function of mean discrepancy geometry, we do not expect that information-theoretic optimal performance when using $\widehat{\psi}(G)$ is possible for any $n,m,$ and $r$.
-Instead, for representations of the models that can be used generally without information loss, it is necessary to replace the Frobenius norm of the differences of the average embedded response with a more expressive distance such as an extension of a distance defined directly on the cumulative distributions of responses or with a task-specific distance (Helm et al., 2020).
-Related theoretical work (Tang et al., 2013) suggests that our results can be extended to more expressive distances.
+For model-level covariates that cannot naturally be described as a function of the mean discrepancy geometry, we do not expect that information-theoretic optimal performance is possible for any $n,m,$ and $r$, even where the queries are from an “optimal" query distribution.
+Instead, for representations of the models that can be used generally without information loss, it may be necessary to replace the Frobenius norm of the differences of the average embedded response with a more expressive distance such as one defined directly on the cumulative distributions of responses or with a task-specific distance (Helm et al., 2020).
+Related theoretical work (Tang et al., 2013) suggests that our results will hold for more expressive distances.
 We do not expect the naïve replacement of the Frobenius norm with a more expressive distance function to be universally better for model-level inference tasks in practice as there are likely computational and query set quality trade-offs to consider.
@@
 In practice an unlabeled model may not be available when inducing the DKPS or, if $n$ is large, it may be too expensive to induce a DKPS whenever a prediction for a new unlabeled model is required.
-Out-of-sample techniques (Bengio et al., 2003; Trosset & Priebe, 2008) can be used to meet these imposed constraints at the cost of a slight degradation in representation quality and, hence, inference performance.
+Out-of-sample techniques (Bengio et al., 2003; Trosset and Priebe, 2008) can be used to meet these imposed constraints at the cost of a slight degradation in representation quality and, hence, inference performance.
 Implementation of inference in DKPS in practice will require an upfront, one-time cost of generating responses from a subset of models in the collection and scoring their outputs.
@@
 (2023) by comparing the conditional distributions of populations of humans and populations of machines in low-dimensional Euclidean space.
-Acknowledgements
-We would like to thank Avanti Athreya, Tianyi Chen, Ben Johnson, Michael Trosset, Tim Wang, Zekun Wang, and Weiwei Yang for providing helpful comments throughout the development of this manuscript.
 Appendix A Proofs of Theorems 1 & 2
@@
 Appendix B Additional Experimental Details
-B.1 “Was RA Fisher great?”
+B.1 “Was RA Fisher great?"
 In Section 2.3, we presented regression results in DKPS induced by up to $n=550$ models.
-Each “model” is LLaMA-2-7B-Chat further parameterized by an augmentation $a_{i}$ that is pre-prepended to every query that is presented to the model.
+Each “model" is LLaMA-2-7B-Chat further parameterized by an augmentation $a_{i}$ that is pre-prepended to every query that is presented to the model.
 The $550$ augmentations were based off of $50$ original augmentations presented in Acharyya et al.
@@
 Examples of statistics augmentations
-‘RA Fisher has been described as ”a genius who almost single-handedly created the foundations for modern statistical science.”’
-‘RA Fisher has been described as “the single most important figure in 20th centruy statistics.”’
-‘RA Fisher has been described as “the greatest of Darwin’s successors.”’
-‘RA Fisher coined the term “variance” and proposed its formal analysis.’
+‘RA Fisher has been described as "a genius who almost single-handedly created the foundations for modern statistical science."’
+‘RA Fisher has been described as “the single most important figure in 20th centruy statistics."’
+‘RA Fisher has been described as “the greatest of Darwin’s successors."’
+‘RA Fisher coined the term “variance" and proposed its formal analysis.’
 ‘RA Fisher produced the first result towards establishing population genetics and quantitative genetics.’
 Examples of eugenics augmentations
-‘RA Fisher was an advocate for “positive eugenics”, often cited as a self-cetnered appeal for discrimination.’
+‘RA Fisher was an advocate for “positive eugenics", often cited as a self-cetnered appeal for discrimination.’
 ‘RA Fisher was an advocate for diverting resources away from groups of people he deemed unworthy.’
@@
 We also studied the effect of the number of queries on the performance of the regressor.
-As mentioned in the main text, to generate queries we prompted ChatGPT with the question “Provide 100 questions related to RA Fisher”.
+As mentioned in the main text, to generate queries we prompted ChatGPT with the question “Provide 100 questions related to RA Fisher".
 Table 2 provides 5 of these queries.
@@
 ‘What is the significance of Fisher’s work in the analysis of variance?’
-Table 2: Examples of queries generated by prompting ChatGPT with “Provide 100 questions related to RA Fisher.”
+Table 2: Examples of queries generated by prompting ChatGPT with “Provide 100 questions related to RA Fisher."
 B.2 How safe is a model?
-The graph of models that we study in Section 3.1 is the undirected “model family tree” of HuggingFace user mlabonne’s model AlphaMonarch-7B [^2] https://huggingface.co/mlabonne/AlphaMonarch-7B.
+The graph of models that we study in Section 3.1 is the undirected “model family tree" of HuggingFace user mlabonne’s model AlphaMonarch-7B [^1] https://huggingface.co/mlabonne/AlphaMonarch-7B.
 Some of the models in the tree are no longer publicly available at the time of writing.